OpenStudy (anonymous):

What is the 19th term of the geometric sequence ? -2,4, -8 , 16 .... Can someone please help? thanks.

OpenStudy (campbell_st):

can you identify the common ratio... what each term is being multiplied by to get the next value

OpenStudy (anonymous):

what do you mean ?

OpenStudy (eric_d):

r=?

OpenStudy (anonymous):

Oh it doesn't say :(

OpenStudy (campbell_st):

-2 x ? = 4 4 x ? = -8 -8 x ? = 16 whats the value of ?

OpenStudy (anonymous):

R= -2 I think.

OpenStudy (eric_d):

4 divide by -2, you're rite

OpenStudy (eric_d):

it's -2

OpenStudy (anonymous):

Ok so now what do I do ?

OpenStudy (anonymous):

and a=-2 to right?

OpenStudy (campbell_st):

thats correct... now the formula to use is \[a_{n} = a \times r^{n - 1}\] from the information a = -2 and n = 19, and you found r = -2 so substitute the, and calcuate the answer

OpenStudy (anonymous):

ok thnk u :)

OpenStudy (campbell_st):

for you to get the right answer you'll need \[a_{19} = -2 \times (-2)^{19 - 1}\]

OpenStudy (eric_d):

-2(-2)^18

OpenStudy (anonymous):

is the answer 990 for the 19th?

OpenStudy (anonymous):

or no

OpenStudy (campbell_st):

try \[a_{19} = -2 \times (-2)^{18}\]

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ok that will be \[a19=-524288\] correct?

OpenStudy (campbell_st):

thats it

OpenStudy (anonymous):

-524288

OpenStudy (anonymous):

thanks so much